1,3,3,3,5,5,5,5,5,7,7,7,7,7,7,7..... what will be its 2016th term
 

Grade:11

1 Answers

RAGHU RAM
6 years ago
1+3*3+5*5+7*7+-----Here 1,3,5,----are odd numbers andnth term of odd numbers is 2n-144²= 1936 and 45²= 2025So the number appears1937th term to 2025th is 45th odd numberi.e 2(45)-1=90-1=89